Shift report 20011110 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Nov 10 2001 - 16:03:23 EST

  • Next message: Apache: "Shift report 20011110 16:00-24:00"

    ------------------------------------------------------------
    Date: Sat Nov 10 16:03:23 2001, Shift: 08:00-16:00
    Supervisor: rw, with: 
    Log book#10 pages 122 - 123
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam from the previous shift was dumped at 10:20. STAR
    took experimental area acess. Beam was preparing to inject 
    at 13:30 and cogged at 14:30, but beam intensity was low for
    PHENIX and dumped after 10 mini. No luck for beam after that.
    
    2.   run      Trigger 6       Setting
         5749      120K         FS 20deg-1/10B  MRS 90deg-2A
         5750       61K          same
         5751       TPC Pedestal run
         5752       DVM
         5773       Pedestals
         5774        2K         FS 30deg-1/15B  MRS 90deg-4A
    
    3. No new problems
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Nov 10 2001 - 16:04:06 EST